# Asynchronous Microphone Array Calibration using Hybrid TDOA Information

### v3

@@ Abstract
Asynchronous microphone array calibration is a prerequisite for most audition robot applications. A popular solution to the above calibration problem is the batch form of Simultaneous Localisation and Mapping (SLAM), using the time difference of arrival measurements between two microphones (TDOA-M), and the robot (which serves as a moving sound source during calibration) odometry information. In this paper, we introduce a new form of measurement for microphone array calibration, i.e. the time difference of arrival between adjacent sound events (TDOA-S) with respect to the microphone channels. We propose to combine TDOA-S and TDOA-M, called hybrid TDOA, together with odometry measurements for bath SLAM-based calibration of asynchronous microphone arrays. Simulation and real-world experiment results consistently show that our method is more independent of microphone number, less sensitive to initial values (when using off-the-shelf algorithms such as Gauss-Newton iterations), and has better calibration accuracy and robustness under various TDOA noises. In addition, the simulation result demonstrates that our method has a lower Cramér-Rao lower bound (CRLB) for microphone parameters. To benefit the community, we open-source our code and data at https://github.com/zcj808/Hybrid-TDOA-Calib.
@@ I INTRODUCTION
Microphone arrays can equip robots with sound source localization and tracking abilities, etc [1, 2, 3, 4, 5, 6]. A prerequisite for realizing the above functionalities is to calibrate the array geometric information accurately [7]. A common approach to the above calibration problem is to utilize the time difference of arrival between microphone pairs (TDOA-M) from a series of sound events to estimate both microphone and sound source locations. Earlier methods require the clock synchronization of all microphones [8, 9]. To overcome the limitation, recent studies, including [10, 11, 12], take into account the initial time offset between microphone channels.
@@
During calibration, one can obtain the relative position measurements between adjacent sound events from the odometer onboard the robot (which acts as a moving sound source) and use them to improve the calibration accuracy. Following the above idea, based on TDOA-M and odometry measurements, an extended Kalman filter-based simultaneous localization and mapping (EKF-SLAM) method is proposed in [13] to estimate microphone positions, time offsets, and sound source positions simultaneously. However, the impact of the other asynchronous factor, i.e. clock drift rate, is not considered. In [14, 15, 16], a batch SLAM-based method [17] is presented to jointly estimate microphone locations, time offsets, clock drift rates, and sound event positions. However, when using off-the-shelf algorithms such as Gauss-Newton iterations, this method requires a good initial value to achieve reasonable calibration results.
@@
    We introduce TDOA-S and present a simple method of extracting TDOA-S in practice. To our best knowledge, this is the first time TDOA-S has been proposed in the literature and used in calibrating robot audition systems. The idea can be generalized to other sensing modalities.
@@
    We propose a batch SLAM-based calibration method utilizing hybrid TDOA information and odometry measurements to jointly estimate the asynchronous microphone array parameters (microphone positions, time offsets, clock drift rates) and sound source positions. Extensive simulations and real-world experiments have been conducted to validate the advantages of the proposed method. The results show that compared to [17], our method is more independent of microphone number, less sensitive to initialization, has higher accuracy and robustness under various TDOA noises, and has lower CRLB for microphone parameter estimates.
@@ II-A1 Derivation
where $c$ is the sound speed and $t_{j}$ is the emitting time of $j$-th sound event. $\tau_{i}$ and $\delta_{i}$ represent the shift and scaling of the temporal frame of $i$-th microphone with respect to (w.r.t.) the absolute temporal frame, respectively. The former is caused by different startup moments in different microphones. The latter is caused by the sampling rate mismatch between the microphone’s actual and absolute sampling rates [18], which can be modeled as a scale constant between a microphone’s temporal frame and the absolute temporal frame. If the mismatch does not exist for $i$-th microphone ($\delta_{i}=0$), the TOA model is the same as the common TOA that only considers time offset [7].
@@ II-A2 Extraction
There are two steps in obtaining TDOA-S, each visualized in Fig. 2. Initially, all audio segments containing a calibration signal in a single-channel microphone are captured for a time window. The rough time delay of adjacent calibration signals ($T_{rough}$) is equal to the difference between the left endpoint of the adjacent windows containing the calibration signal (Fig. 2a). Next, we align the adjacent windows and perform GCC-PHAT [19] to obtain the precise delay ($T_{pre}$) (see Fig. 2b). Finally, one combines the rough delay and precise delay to obtain the overall delay ($T_{rough}+T_{pre}$), which is equal to the difference between two consecutive moments of arrivals.
@@ II-B1 Hybrid TDOA Measurements
where $\tau_{i,1}=\tau_{i}-\tau_{1}$, $\delta_{i,1}=\delta_{i}-\delta_{1}$ ($i>1$). Assume $t_{1}=0$ without loss of generality, for $j>1$, $t_{j}=t_{j}-t_{1}=\sum^{j}_{k=2}\Delta t_{k-1}$. The TDOA-M formula (6) is equivalent to the TDOA formula in [16]. Hence, without noise, the total hybrid TDOA measurements are
@@
where $\mathbf{T}^{S}=[\mathbf{T}^{S}_{1},\mathbf{T}^{S}_{2},...,\mathbf{T}^{S}_{N}]^{T}$, $\mathbf{T}^{S}_{i}=[T^{S}_{i,1},T^{S}_{i,2},...,T^{S}_{i,K-1}]^{T}$ and $\mathbf{T}^{M}=[\mathbf{T}^{M}_{1},\mathbf{T}^{M}_{2},...,\mathbf{T}^{M}_{K}]^{T}$, where $\mathbf{T}^{M}_{j}=[T^{M}_{2,j},T^{M}_{3,j},...,T^{M}_{N,j}]^{T}$.
@@
II-B2 Nonlinear Least Squares solved by GN method
@@ III-1 Setup
In “True $\mathbf{x}_{mic}$”, “random” means microphone locations are randomly generated in the corresponding trajectory space and $|\tau_{i,1}|\leq 0.1s$, $|\delta_{i}|\leq 10^{-4}s$. In “Initial $\mathbf{x}_{mic}$”, “random” means both microphone and sound event locations are randomly generated in the corresponding trajectory space, and asynchronous parameters are set to be zero. $\sigma_{init}$ are standard deviations (SDs) of zero-mean Gaussian noises added into the true positions as the initial values of both microphone and sound event locations. In trajectory 1, $\sigma_{init}=0m,1m,2m,3m$ and in trajectory both 2 and 3, $\sigma_{init}=0m,2m,4m,6m$. Simulation under different numbers of microphones (Part A), various initial value noises (Part B), and several TDOA noises (Part C/D) repeat 200 times in each trajectory and the results of three trajectories are combined to analyze.
@@ III-3 Results
It can be observed from Fig. 3a-c that as the number of microphones changes, the calibration performance for microphone parameters of our method remains basically unchanged. However, the performance of [15] shows significant changes and approaches that of our method as the number increases. Fig. 3d-f shows that the estimation performance for microphone parameters of our method remains unchanged under different initial value noises. However, microphone parameters estimated by [15] exhibit an increase in estimation error as the initial values noise increases. In Fig. 3g-i, we can observe that our method has better accuracy and robustness in estimating microphone parameters under three levels of TDOA noises, as we have lower median and IQR values for each box. Table I confirms that our method estimates the CRLB for microphone parameters to be smaller under various TDOA noises.
@@ IV-4 Results
Figure 5: Box plot of results in the real-world experiment: microphone location estimation errors under various microphone numbers (a), i.e. 4, 6, 8, and 10, initial values noise SDs (b), i.e. 0m, 0.5m, 1m, and 2m, and five cases of TDOA noises (c).
@@ VI CONCLUSIONS
This paper is concerned with asynchronous microphone array calibration using batch SLAM. More specifically, we have introduced a new type of measurement, i.e. TDOA-S, for the above calibration problem. We have presented a simple procedure to extract TDOA—S measurements from raw audio information. We proposed to use hybrid TDOA information (both TDOA-S and TDOA-M) for calibrating asynchronous microphone arrays. Both simulation and experiment results show that our proposed method leads to more desirable calibration results than the case using only TDOA-M information. In particular, the proposed method is more independent of the number of microphones, has lower sensitivity to initial values, and has higher accuracy and robustness under various TDOA noises. The focus of our current and future work is to explore the relationship between calibration observability (eigenvalues of the Fisher information matrix) and sound source trajectory configuration and generalize the ideas in this paper to the problem of calibrating multiple microphone arrays.
@@ VII-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
$\tilde{t}^{S}_{i,j}=t^{S}_{i,j}-\frac{||\mathbf{x}_{i}-\mathbf{s}_{j+1}||-||\mathbf{x}_{i}-\mathbf{s}_{j}||}{c}-\Delta t_{j}=\delta_{i}\Delta t_{j}+w^{S}_{i,j}.$

### v4

@@ Abstract
Asynchronous microphone array calibration is a prerequisite for many audition robot applications. A popular solution to the above calibration problem is the batch form of Simultaneous Localisation and Mapping (SLAM), using the time difference of arrival measurements between two microphones (TDOA-M), and the robot (which serves as a moving sound source during calibration) odometry information. In this paper, we introduce a new form of measurement for microphone array calibration, i.e. the time difference of arrival between adjacent sound events (TDOA-S) with respect to the microphone channels. We propose to use TDOA-S and TDOA-M, called hybrid TDOA, together with odometry measurements for bath SLAM-based calibration of asynchronous microphone arrays. Extensive simulation and real-world experiments show that our method is more independent of microphone number, less sensitive to initial values (when using off-the-shelf algorithms such as Gauss-Newton iterations), and has better calibration accuracy and robustness under various TDOA noises. Simulation results also demonstrate that our method has a lower Cramér-Rao lower bound (CRLB) for microphone parameters. To benefit the community, we open-source our code and data at https://github.com/AISLAB-sustech/Hybrid-TDOA-Calib.
@@ I INTRODUCTION
Microphone arrays can equip robots with sound source localization and tracking abilities, etc [1, 2, 3, 4, 5]. A prerequisite for realizing the above functionalities is to calibrate the array geometric information accurately [6]. A common approach to the above calibration problem is to utilize the time difference of arrival between microphone pairs (TDOA-M) from a series of sound events to estimate both microphone and sound source locations. Earlier methods require the clock synchronization of all microphones [7, 8]. To overcome the limitation, recent studies, including [9, 10, 11], take into account the initial time offset between microphone channels.
@@
During calibration, one can obtain the relative position measurements between adjacent sound events from the odometer onboard the robot (which acts as a moving sound source) and use them to improve the calibration accuracy. Following the above idea, based on TDOA-M and odometry measurements, an extended Kalman filter-based simultaneous localization and mapping (EKF-SLAM) method is proposed in [12] to estimate microphone positions, time offsets, and sound source positions simultaneously. However, the impact of the other asynchronous factor, i.e. clock drift rate, is not considered. In [13, 14], a batch SLAM-based method (see, e.g., [15]) is presented to jointly estimate microphone locations, time offsets, clock drift rates, and sound event positions. In our recent work, the framework in [13, 14] has also been generalized to multiple microphone arrays [16] and refined by the data compression method [17].
@@
    We introduce TDOA-S and present a simple method to extract TDOA-S from raw audio data. To our best knowledge, this is the first time TDOA-S has been proposed in the literature and used in calibrating robot audition systems. The idea can be useful for other sensing modalities.
@@
    We propose a batch SLAM-based calibration method utilizing hybrid TDOA information and odometry measurements to jointly estimate the asynchronous microphone array parameters (microphone positions, time offsets, clock drift rates) and sound source positions. Extensive simulations and real-world experiments show that compared to [13], our method is more independent of microphone number, less sensitive to initialization, has higher accuracy and robustness under various TDOA noises, and has lower CRLB for microphone parameter estimates.
@@ II-A1 Derivation
where $c$ is the known sound speed and $t_{j}$ is the emitting time of $j$-th sound event. $\tau_{i}$ and $\delta_{i}$ represent the shift and scaling of the temporal frame of $i$-th microphone with respect to (w.r.t.) the absolute temporal frame, respectively. The former is caused by different startup moments in different microphones. The latter is caused by the sampling rate mismatch between the microphone’s actual and absolute sampling rates [18], which can be modeled as a scale constant between a microphone’s temporal frame and the absolute temporal frame. If the mismatch does not exist for $i$-th microphone ($\delta_{i}=0$), the TOA model is the same as the common TOA that only considers time offset [6].
@@ II-A2 Extraction
There are two steps in obtaining TDOA-S, each visualized in Fig. 2. Initially, short-time energy [19] is used to obtain the rough left endpoint of each calibration signal in a single-channel microphone. The rough time delay of adjacent calibration signals ($T_{rough}$) is equal to the difference between the corresponding rough left endpoints (Fig. 2a). Next, each window containing a calibration signal is extracted based on the signal length and rough left endpoint, and we align the adjacent windows and perform GCC-PHAT [20] to obtain the precise delay ($T_{pre}$) (see Fig. 2b). Finally, one combines the rough delay and precise delay to obtain the overall delay ($T_{rough}+T_{pre}$), which is equal to the difference between two consecutive moments of arrivals.
@@ II-B1 Hybrid TDOA Measurements
where $\delta_{i,1}=\delta_{i}-\delta_{1}$ ($i>1$). Assume $t_{1}=0$ without loss of generality, for $j>1$, $t_{j}=t_{j}-t_{1}=\sum^{j}_{k=2}\Delta t_{k-1}$. The TDOA-M formula (6) is equivalent to the TDOA formula in [13]. Hence, without noise, the total hybrid TDOA measurements are
@@
II-B2 Nonlinear Least Squares solved by Gauss-Newton
@@ III-1 Setup
In “True $\mathbf{x}_{mic}$”, “random” means microphone locations are randomly generated in the corresponding trajectory space and $|\tau_{i,1}|\leq 0.1s$, $|\delta_{i}|\leq 10^{-4}s$. In “Initial $\mathbf{x}$”, “random” means both microphone and sound event locations are randomly generated in the corresponding trajectory space, and asynchronous parameters are set to be zero. $\sigma_{init}$ are standard deviations (SDs) of zero-mean Gaussian noises added into the true positions as the initial values of both microphone and sound event locations. In trajectory 1, $\sigma_{init}=0m,1m,2m,3m$ and in trajectory both 2 and 3, $\sigma_{init}=0m,2m,4m,6m$. Simulation under different numbers of microphones (Part A), various initial value noises (Part B), and several TDOA noises (Part C/D) repeat 200 times in each trajectory and the results of three trajectories are combined to analyze.
@@ III-3 Results
It can be observed from Fig. 3a-c that as the number of microphones changes, the calibration performance for microphone parameters of our method remains basically unchanged. However, the performance of [13] shows significant changes and approaches that of our method as the number increases. Fig. 3d-f shows that the estimation performance for microphone parameters of our method remains unchanged under different initial value noises. However, microphone parameters estimated by [13] exhibit an increase in estimation error as the initial values noise increases. In Fig. 3g-i, we can observe that our method has better accuracy and robustness in estimating microphone parameters under three levels of TDOA noises, as we have lower median and interquartile range (IQR) values for each box. Table II confirms that our method estimates the CRLB for microphone parameters to be smaller under various TDOA noises.
@@ IV-4 Results
Figure 5: Box plot of results in real-world experiment: microphone location estimation errors under various microphone numbers (a), i.e. 4, 6, 8, and 10, initial values noise SDs (b), i.e. 0m, 0.5m, 1m, and 2m, and five cases of TDOA noises (c).
@@ VI CONCLUSIONS
This paper is concerned with asynchronous microphone array calibration using batch SLAM. More specifically, we have introduced a new type of measurement, i.e. TDOA-S, for the above calibration problem. We have presented a simple procedure to extract TDOA-S measurements from raw audio information. We proposed to use hybrid TDOA information (both TDOA-S and TDOA-M) for calibrating asynchronous microphone arrays. Both simulation and experiment results show that our proposed method leads to improved calibration results compared to the case using only TDOA-M information. In particular, the proposed method is more independent of the number of microphones, has lower sensitivity to initial values, and has higher accuracy and robustness under various TDOA noises. The focus of our current and future work is to explore the relationship between calibration observability and sound source trajectory configuration and generalize the ideas in this paper to the problem of calibrating multiple microphone arrays.
@@ VII-B1 Computation of $\tilde{\sigma}^{S}_{tdoa}$
$\tilde{t}^{S}_{i,j}$ $=t^{S}_{i,j}-\frac{||\mathbf{x}_{i}-\mathbf{s}_{j+1}||-||\mathbf{x}_{i}-\mathbf{s}_{j}||}{c}-\Delta t_{j}$
$=\delta_{i}\Delta t_{j}+w^{S}_{i,j}.$